ABSTRACT. Some recent results are described concerning the characterization and computation of minimal norm elements from a subset of a Hilbert space. This subset is the intersection of a closed convex cone and a closed linear variety. Many important problems in statistics, mathematical modeling, and curve and surface fitting fall in this category.